# Neural Wave Functions for Superfluids

### v3

@@ III Antisymmetrized Geminal Power Wave Function
The FermiNet and other Ansätze that generalize the idea of a Slater determinant or expand the ground state as a linear combination of Slater determinants give very accurate results for many molecules and solids, but may still fail to capture strong two-particle correlations in superfluids. An alternative starting point, which is better at capturing two-particle correlations, is the antisymmetric geminal power (AGP) wave function [40, 51, 41, 42]. This uses an antisymmetrized product of two-particle functions known as pairing orbitals or geminals instead of an antisymmetrized product of single-particle orbitals.
@@ III.1 AGP Singlet Wave Functions
Like all spin-assigned wave functions, it depends on position coordinates only. For convenience, we have changed the particle labeling scheme: $i$ and $j$ both run from $1$ to $p$ and arrow superscripts have been added to distinguish up-spin from down-spin particles. Note that the AGPs wave function coincides with the BCS wavefunction projected onto a fixed particle number subspace (see Appendix) [39]. It is therefore suitable for describing singlet-paired systems, including $s$-wave superfluids.
@@ III.2 AGPs with Unpaired States
where $\varphi(\mathbf{r}^{\uparrow}_{i},\mathbf{r}^{\downarrow}_{j})$ is the singlet pairing function and $\phi^{\sigma_{i}}_{i}(\mathbf{r}^{\sigma_{i}}_{j})$ are the single-particle orbitals. For the UFG considered in this paper, we only need the case where $u=1$ and $d=0$ or vice versa. This represents a fully paired $2p$-particle system to which one particle has been added. The extra particle is placed in a single-particle orbital.
@@ IV AGP Singlet FermiNet
Having discussed the form of the AGPs wave function, we now discuss how it can be implemented using FermiNet. In the original FermiNet architecture, the outputs of the one-electron stream are used to build FermiNet orbitals $\phi^{k\alpha}_{i}(\mathbf{r}^{\alpha}_{j};\{\mathbf{r}^{\alpha}_{/j}\};\{\mathbf{r}^{\bar{\alpha}}\})$. The full many-fermion wave function is a weighted sum of terms, each of which is the product of one up-spin and one down-spin determinant of the FermiNet orbital matrices, as shown in Eq. (11).
To build a pairing orbital, one can make use of the outputs $\mathbf{h}^{L\alpha}_{i}$ from the last layer $L$ of the one-electron stream. Instead of using these outputs to build FermiNet orbitals, as in Eq. (9), they can be used to build FermiNet pairing orbitals as follows:
$\varphi^{k}(\mathbf{r}^{\alpha}_{i},\mathbf{r}^{\bar{\alpha}}_{j};\{\mathbf{r}^{\alpha}_{/i}\};\{\mathbf{r}^{\bar{\alpha}}_{/j}\})\\ =[\mathbf{w}^{k}\cdot(\mathbf{h}^{L\alpha}_{i}\odot\mathbf{h}^{L\bar{\alpha}}_{j})]\chi^{k}(\mathbf{r}^{\alpha}_{i})\chi^{k}(\mathbf{r}^{\bar{\alpha}}_{j}),$ (17)
where $\chi(\mathbf{r})$ are the envelope functions, $\mathbf{w}^{k}$ are vectors and $\odot$ denotes the element-wise product. This procedure generates a pairing function between particles $\mathbf{r}^{\alpha}_{i}$ and $\mathbf{r}^{\bar{\alpha}}_{j}$, retaining the permutation invariant property possessed by FermiNet orbitals. Depending on the number of FermiNet geminals generated, the AGPs can be written as one or a weighted sum of multiple determinants of the FermiNet geminals
$\Psi^{\text{AGPs}}_{\text{FermiNet}}\left(\mathbf{r}^{\uparrow}_{1},\dots,\mathbf{r}^{\downarrow}_{N^{\downarrow}}\right)$
$=\sum_{k}\omega_{k}\det[\varphi^{k}(\mathbf{r}^{\alpha}_{i},\mathbf{r}^{\bar{\alpha}}_{j};\{\mathbf{r}^{\alpha}_{/i}\};\{\mathbf{r}^{\bar{\alpha}}_{/j}\})].$ (18)
This is analogous to a weighted sum of single-determinant AGPs wave functions of the type defined in Eq. (15), but the replacement of the two-particle geminals by FermiNet geminals that depend on the positions of all of the particles makes it much more general.
@@
$+\mathbf{w}^{k}_{2}\cdot(\mathbf{h}^{L\alpha\bar{\alpha}}_{ij}\odot\mathbf{h}^{L\bar{\alpha}\alpha}_{ji})]\chi^{k}(\mathbf{r}^{\alpha}_{i})\chi^{k}(\mathbf{r}^{\bar{\alpha}}_{j}),$ (19)
@@
Note that Eqs. (17) and (19) are two possible ways of building a pairing function. There are many others ways and they are all valid as long as the appropriate symmetries are preserved. An alternative method is given by Xie et al. [52].
The benefit of building AGPs-like wave functions using the FermiNet is that the pairing function $\varphi(\mathbf{r}^{\alpha}_{i},\mathbf{r}^{\bar{\alpha}}_{j};\{\mathbf{r}^{\alpha}_{/i}\};\{\mathbf{r}^{\bar{\alpha}}_{/j}\})$ now depends not only on $\mathbf{r}_{i}^{\alpha}$ and $\mathbf{r}_{j}^{\bar{\alpha}}$ but also on the positions of the other particles in the system 222 Note that the idea of an AGPs/BCS FermiNet is very similar to the previous work by Luo and Clark [30], where they used the neural network backflow (NNB) wave function implemented on top of a Bogoliubov-de Gennes/BCS wave function to study lattice systems. . Correlations between the singlet pair and the other particles can thus be captured. In a similar way, the original unpaired FermiNet replaced Hartree-Fock-like single-particle orbitals $\phi_{i}^{\alpha}(\mathbf{r}_{j}^{\alpha})$ by FermiNet orbitals $\phi^{\alpha}_{i}(\mathbf{r}^{\alpha}_{j};\{\mathbf{r}^{\alpha}_{/j}\};\{\mathbf{r}^{\bar{\alpha}}\})$, helping to capture correlations between the particle at $\mathbf{r}_{j}^{\alpha}$ and all other particles.
IV.1 Relations between the original FermiNet and the AGPs FermiNet
Next, we clarify the relation between the original FermiNet and the AGPs FermiNet, showing that the AGPs FermiNet is the more general of the two. A FermiNet geminal with a two-particle stream term is even more general than a FermiNet geminal without, so it is sufficient for this purpose to omit the two-particle stream term. We also neglect the envelope functions. Including them circumvents numerical difficulties in finite systems and can speed up the network optimization, but does not affect the generality of the Ansatz.
Let us first define a pairing function in the following way:
@@ IV.1 Relations between the original FermiNet and the AGPs FermiNet
We choose to write the pairing function in the form of Eq. (20) only because this makes it easier to relate to FermiNet orbitals. Since Eqs. (20) and (21) are equivalent, the choice does not affect the conclusions of the argument. In the rest of this section, for the sake of simplicity, we omit the sets $\{\mathbf{r}_{/i}^{\uparrow}\}$ and $\{\mathbf{r}_{/j}^{\downarrow}\}$ from the arguments of FermiNet pairing functions and orbitals.
@@
where $U\in\mathbb{R}^{n_{L}\times n_{L}}$ and $V\in\mathbb{R}^{n_{L}\times n_{L}}$ are orthogonal matrices and $n_{L}$ is the size of the vectors $\mathbf{h}_{i}^{L\alpha}$ output by the final layer $L$ of the one-electron stream. This is also known as the number of hidden units in layer $L$. The pairing function in Eq. (20) becomes
@@
$=\sum_{\gamma}\sigma_{\gamma}\left(\sum_{k}U_{\gamma k}h^{L\uparrow(k)}_{i}\right)\left(\sum_{k}V_{\gamma l}h^{L\uparrow(l)}_{j}\right)$ (24)
@@
Given the universal approximation theorem [44], and the fact that every layer of the network contains an arbitrary linear transformation, it is reasonable to assume that the functions $\mathbf{h}^{L\alpha}_{i}$ and $O\mathbf{h}^{L\alpha}_{i}$, where $O$ is $U$ or $V$, have the same variational freedom and information content. In other words, we assume that any network capable of representing $\mathbf{h}^{L\alpha}_{i}$ can also represent $O\mathbf{h}^{L\alpha}_{i}$, since this is merely a rotation of the vectors $\mathbf{h}^{L\alpha}_{i}$ in the last layer. We thus define $\tilde{\mathbf{h}}^{L\uparrow}_{i}=U\mathbf{h}^{L\uparrow}_{i}$ and $\tilde{\mathbf{h}}^{L\downarrow}_{i}=V\mathbf{h}^{L\downarrow}_{i}$, such that the pairing function becomes
@@
To relate the AGPs FermiNet and the original FermiNet, we expand an AGPs determinant constructed using the pairing function from Eq. (20) as a sum of block-diagonal determinants of FermiNet orbitals. It will be sufficient to consider matrices $W\in\mathbb{R}^{n_{L}\times n_{L}}$ of rank $M$, with $p\leq M\leq n_{L}$. We can decompose any such matrix using rank factorization,
@@
where the last line defines the functions $\phi_{\gamma}^{\uparrow}$ and $\phi_{\gamma}^{\downarrow}$. In the case when $M=p$, where $p=N/2$ is the number of pairs in the system, the determinant of the pairing function can be written as a block-diagonal determinant of FermiNet orbitals:
@@
where $[\phi^{\alpha}]_{ij}=\phi^{\alpha}_{i}(\mathbf{r}^{\alpha}_{j})$ are matrices in $\mathbb{R}^{M\times p}$ with $M=p$. The product of two $p\times p$ determinants can be written as the determinant of a single $2p\times 2p$ matrix, with the $p\times p$ spin-up and spin-down blocks on the diagonal. Therefore, any AGPs wave function constructed using the pairing function from Eq. (20) with a rank-$p$ matrix $W_{kl}$ is equivalent to a $2p\times 2p$ block-diagonal determinant of conventional FermiNet orbitals. The equivalence is already well known [50] for APGs wave functions constructed using conventional two-particle orbitals.
@@
where the sum is over all $\binom{M}{p}$ distinct choices of $p$ rows from the two $M\times p$ matrices $\phi^{\uparrow}_{\gamma}(\mathbf{r}^{\uparrow}_{i})$ and $\phi^{\downarrow}_{\gamma}(\mathbf{r}^{\downarrow}_{i})$. The products of the determinants of the two $p\times p$ matrices associated with each such choice are summed to reproduce the APGs. This is similar to the block-diagonal multiple determinant expansion of FermiNet orbitals without weights given by Eq. (11) and in the original FermiNet paper [1] 333 To get a rough estimate of the number of terms in the sum, we can take the number of hidden units in the one-electron stream to be $M=n_{L}\sim 200$ and the number of pairs to be $p\sim 20$, this number is approximately $10^{27}$. .
Note that the intermediate layers, i.e., the one and two-electron streams, are identical in the original FermiNet and the AGPs FermiNet. The only modifications are made at the orbital shaping layer, or, equivalently, the method of antisymmetrization has changed. Thus, the representational power of the intermediate layers of the AGPs FermiNet remain the same as for the original FermiNet. Thus, it must be the method of antisymmetrization that limits the performance of the original FermiNet when applied to the UFG.
We have shown that a single APGs determinant constructed using the pairing function from Eq. (20) with a matrix $W_{kl}$ of rank greater than $p$ contains multiple block-diagonal determinants of FermiNet orbitals. If the rank of $W_{kl}$ is equal to $p$, the AGPs is equivalent to a single block-diagonal FermiNet determinant. Conversely, any single-determinant FermiNet wave function can be written as an AGPs of rank $p$. Therefore, the AGPs FermiNet provide a more powerful Ansatz with fewer variational parameters than the original FermiNet, since the former contains the latter.
It is worth mentioning another advantage of using the FermiNet to build geminals. By generating more sets of independent parameters $\mathbf{w}^{k}_{i}$ in Eq. (19), one can easily construct an arbitrary number $N_{\det}$ of FermiNet geminals $\varphi^{k}(\mathbf{r}^{\alpha}_{i},\mathbf{r}^{\bar{\alpha}}_{j};\{\mathbf{r}^{\alpha}_{/i}\};\{\mathbf{r}^{\bar{\alpha}}_{/j}\})$ with $k\in\{1,2,\dots,N_{\det}\}$, all without the use of a basis set. It allows one to use weighted sums of AGPs determinants as trial wave functions, similar to the weighted sum of conventional FermiNet determinants seen in Eq. (11).

### v4

@@ III Antisymmetrized Geminal Power Wave Function
The FermiNet and other Ansätze that expand the ground state as a linear combination of Slater determinants give very accurate results for many molecules and solids, but may still fail to capture strong two-particle correlations in superfluids. An alternative starting point, which is better at capturing two-particle correlations, is the antisymmetric geminal power (AGP) wave function [40, 52, 41, 42]. This uses an antisymmetrized product of two-particle functions known as pairing orbitals or geminals instead of an antisymmetrized product of single-particle orbitals.
@@ III.1 AGP Singlet Wave Functions
Like all spin-assigned wave functions, it depends on position coordinates only. For convenience, we have changed the particle labeling scheme: $i$ and $j$ now run from $1$ to $p$ and arrow superscripts have been added to distinguish up-spin from down-spin particles. Note that the AGPs wave function coincides with the BCS wave function projected onto a fixed particle number subspace (see Ref. 39 and Appendix A.1). It is therefore suitable for describing singlet-paired systems, including $s$-wave superfluids.
@@ III.2 AGPs with Unpaired States
where $\varphi(\mathbf{r}^{\uparrow}_{i},\mathbf{r}^{\downarrow}_{j})$ is an arbitrary singlet pairing function and $\phi^{\sigma_{i}}_{i}(\mathbf{r}^{\sigma_{i}}_{j})$ are arbitrary single-particle functions. For the UFG considered in this paper, we only need the case where $u=1$ and $d=0$ or vice versa. This represents a fully paired $2p$-particle system to which one particle has been added.
@@ IV AGP Singlet FermiNet
Having discussed the form of the AGPs wave function, we now discuss how it can be implemented using FermiNet. In the original Slater FermiNet architecture, the outputs of the one-electron stream are used to build FermiNet orbitals $\phi^{k\alpha}_{i}(\mathbf{r}^{\alpha}_{j};\{\mathbf{r}^{\alpha}_{/j}\};\{\mathbf{r}^{\bar{\alpha}}\})$. The full many-particle wave function is a weighted sum of terms, each of which is the product of one up-spin and one down-spin determinant of the FermiNet orbital matrices, as shown in Eq. (11).
To build a many-particle pairing function using the neural-network part of FermiNet, one can make use of its outputs $\mathbf{h}^{L\alpha}_{i}$ from the last layer $L$ of the one-electron stream. Instead of using these outputs to build FermiNet orbitals, as in Eq. (9), they can be used to build FermiNet pairing orbitals, also known as FermiNet geminals:
$\varphi^{k}(\mathbf{r}^{\alpha}_{i},\mathbf{r}^{\bar{\alpha}}_{j};\{\mathbf{r}^{\alpha}_{/i}\};\{\mathbf{r}^{\bar{\alpha}}_{/j}\})\\ =[\mathbf{w}^{k}\cdot(\mathbf{h}^{L\alpha}_{i}\odot\mathbf{h}^{L\bar{\alpha}}_{j})+g^{k}]\chi^{k}(\mathbf{r}^{\alpha}_{i})\chi^{k}(\mathbf{r}^{\bar{\alpha}}_{j}),$ (17)
where $\chi^{k}(\mathbf{r})$ are the envelope functions, $\mathbf{w}^{k}$ are vectors, $g^{k}$ a scalar, and $\odot$ denotes the element-wise product. Note that the same FermiNet geminal is used for all pairs of particles, so the envelope functions in Eq. (17) do not require the particle and spin indices that appear in the envelope functions of the FermiNet orbitals defined in Eq. (9). This construction generates a many-particle pairing function between particles $\mathbf{r}^{\alpha}_{i}$ and $\mathbf{r}^{\bar{\alpha}}_{j}$, retaining the permutation invariant property possessed by FermiNet orbitals. Depending on the number of FermiNet geminals generated, the wave function can be written as one or a weighted sum of multiple determinants of FermiNet geminals,
$\Psi^{D}_{\text{AGPs FermiNet}}\left(\mathbf{r}^{\uparrow}_{1},\dots,\mathbf{r}^{\downarrow}_{N^{\downarrow}}\right)$
$=\sum^{D}_{k}\omega_{k}\det[\varphi^{k}(\mathbf{r}^{\alpha}_{i},\mathbf{r}^{\bar{\alpha}}_{j};\{\mathbf{r}^{\alpha}_{/i}\};\{\mathbf{r}^{\bar{\alpha}}_{/j}\})],$ (18)
where the superscript $D$ in $\Psi^{D}_{\text{AGPs FermiNet}}$ specifies the number of determinants (and thus FermiNet geminals) appearing in the linear combination that makes up the wave function. This is analogous to a weighted sum of conventional single-determinant AGPs wave functions of the type defined in Eq. (15), but the replacement of the two-particle pairing orbitals by FermiNet geminals that depend on the positions of all the particles makes it much more general.
@@
$+\mathbf{w}^{k}_{2}\cdot(\mathbf{h}^{L\alpha\bar{\alpha}}_{ij}\odot\mathbf{h}^{L\bar{\alpha}\alpha}_{ji})+g^{k}]\chi^{k}(\mathbf{r}^{\alpha}_{i})\chi^{k}(\mathbf{r}^{\bar{\alpha}}_{j}),$ (19)
@@
Note that Eqs. (17) and (19) are two possible ways of building a many-particle pairing function. There are many others ways and they are all valid as long as the appropriate symmetries are preserved. An alternative method is given by Xie et al. [53].
The benefit of building AGPs-like wave functions using the FermiNet is that the many-particle pairing function $\varphi(\mathbf{r}^{\alpha}_{i},\mathbf{r}^{\bar{\alpha}}_{j};\{\mathbf{r}^{\alpha}_{/i}\};\{\mathbf{r}^{\bar{\alpha}}_{/j}\})$ now depends not only on $\mathbf{r}_{i}^{\alpha}$ and $\mathbf{r}_{j}^{\bar{\alpha}}$ but also on the positions of the other particles in the system 333 Note that the idea of an AGPs/BCS FermiNet is very similar to the previous work by Luo and Clark [30], where they used the neural network backflow (NNB) wave function implemented on top of a Bogoliubov-de Gennes/BCS wave function to study lattice systems. . Correlations between the singlet pair and the other particles can thus be captured. In a similar way, the original Slater FermiNet replaced Hartree-Fock-like single-particle orbitals $\phi_{i}^{\alpha}(\mathbf{r}_{j}^{\alpha})$ by many-particle orbitals (FermiNet orbitals) $\phi^{\alpha}_{i}(\mathbf{r}^{\alpha}_{j};\{\mathbf{r}^{\alpha}_{/j}\};\{\mathbf{r}^{\bar{\alpha}}\})$, helping to capture correlations between the particle at $\mathbf{r}_{j}^{\alpha}$ and all other particles.
IV.1 Relations between the Slater FermiNet and the AGPs FermiNet
Next, we clarify the relation between the Slater FermiNet with block-diagonal determinants and the AGPs FermiNet, showing that the AGPs FermiNet is the more general of the two. A FermiNet geminal with a two-particle stream term is even more general than a FermiNet geminal without, so it is sufficient for this purpose to omit the two-particle stream term. We also neglect the envelope functions and the bias term, $g^{k}$, which is set to 0 in all results presented here. The use of envelope functions circumvents numerical difficulties in finite systems and can speed up the network optimization, but does not affect the generality of the Ansatz.
Let us first define a many-particle pairing function in the following way:
@@ IV.1 Relations between the Slater FermiNet and the AGPs FermiNet
We choose to write the many-particle pairing function in the form of Eq. (20) only because this makes it easier to relate to FermiNet orbitals. Since Eqs. (20) and (21) are equivalent, the choice does not affect the conclusions of the argument. In the rest of this section, for the sake of simplicity, we omit the sets $\{\mathbf{r}_{/i}^{\uparrow}\}$ and $\{\mathbf{r}_{/j}^{\downarrow}\}$ from the arguments of the many-particle pairing functions and orbitals.
@@
where $U\in\mathbb{R}^{n_{L}\times n_{L}}$ and $V\in\mathbb{R}^{n_{L}\times n_{L}}$ are orthogonal matrices and $n_{L}$ is the size of the vectors $\mathbf{h}_{i}^{L\alpha}$ output by the final layer $L$ of the one-electron stream. This is also known as the number of hidden units in layer $L$. The many-particle pairing function in Eq. (20) becomes
@@
$=\sum_{\gamma}\sigma_{\gamma}\left(\sum_{k}U_{\gamma k}h^{L\uparrow(k)}_{i}\right)\left(\sum_{l}V_{\gamma l}h^{L\uparrow(l)}_{j}\right)$ (24)
@@
Given the universal approximation theorem [44], and the fact that every layer of the network contains an arbitrary linear transformation, it is reasonable to assume that the functions $\mathbf{h}^{L\alpha}_{i}$ and $O\mathbf{h}^{L\alpha}_{i}$, where $O$ is $U$ or $V$, have the same variational freedom and information content. In other words, we assume that any network capable of representing $\mathbf{h}^{L\alpha}_{i}$ can also represent $O\mathbf{h}^{L\alpha}_{i}$, since this is merely a rotation of the vectors $\mathbf{h}^{L\alpha}_{i}$ in the last layer. We thus define $\tilde{\mathbf{h}}^{L\uparrow}_{i}=U\mathbf{h}^{L\uparrow}_{i}$ and $\tilde{\mathbf{h}}^{L\downarrow}_{i}=V\mathbf{h}^{L\downarrow}_{i}$, such that the many-particle pairing function becomes
@@
To relate the AGPs FermiNet and the Slater FermiNet, we expand an AGPs determinant constructed using the many-particle pairing function from Eq. (20) as a sum of block-diagonal determinants of FermiNet orbitals. It will be sufficient to consider matrices $W\in\mathbb{R}^{n_{L}\times n_{L}}$ of rank $M$, with $p\leq M\leq n_{L}$. We can decompose any such matrix using rank factorization,
@@
where the last line defines the functions $\phi_{\gamma}^{\uparrow}$ and $\phi_{\gamma}^{\downarrow}$. In the case when $M=p$, where $p=N/2$ is the number of pairs in the system, the determinant of the many-particle pairing function can be written as a block-diagonal determinant of FermiNet orbitals:
@@
where $[\phi^{\alpha}]_{ij}=\phi^{\alpha}_{i}(\mathbf{r}^{\alpha}_{j})$ are matrices in $\mathbb{R}^{M\times p}$ with $M=p$. The product of two $p\times p$ determinants can be written as the determinant of a single $2p\times 2p$ matrix, with the $p\times p$ spin-up and spin-down blocks on the diagonal. Therefore, a single-geminal AGPs FermiNet wave function constructed using the many-particle pairing function from Eq. (20) with a rank-$p$ matrix $W_{kl}$ is equivalent to a $2p\times 2p$ block-diagonal determinant of FermiNet orbitals. The equivalence is already well known [51] for AGPs wave functions constructed using conventional two-particle orbitals.
@@
where the sum is over all $\binom{M}{p}$ distinct choices of $p$ rows from the two $M\times p$ matrices $\phi^{\uparrow}_{\gamma}(\mathbf{r}^{\uparrow}_{i})$ and $\phi^{\downarrow}_{\gamma}(\mathbf{r}^{\downarrow}_{i})$. The products of the determinants of the two $p\times p$ matrices associated with each such choice are summed to reproduce the AGPs. This is similar to the linear combination of multiple block-diagonal-determinants of FermiNet orbitals without weights given by Eq. (11) and in the original FermiNet paper [1] 444 To get a rough estimate of the number of terms in the sum, we can take the number of hidden units in the one-electron stream to be $M=n_{L}\sim 200$ and the number of pairs to be $p\sim 20$, this number is approximately $10^{27}$. .
Note that the intermediate layers, i.e., the one and two-electron streams, are identical in the Slater FermiNet and the AGPs FermiNet. The only modifications are made at the orbital shaping layer, or, equivalently, the method of antisymmetrization has changed. Thus, the representational power of the intermediate layers of the AGPs FermiNet remains the same as for the Slater FermiNet. Thus, it must be the method of antisymmetrization that limits the performance of the Slater FermiNet when applied to the UFG.
We have shown that a single AGPs determinant constructed using the many-particle pairing function from Eq. (20) with a matrix $W_{kl}$ of rank greater than $p$ contains multiple block-diagonal determinants of FermiNet orbitals. If the rank of $W_{kl}$ is equal to $p$, the AGPs is equivalent to a single block-diagonal FermiNet determinant. Conversely, any single-determinant FermiNet wave function can be written as an AGPs of rank $p$. Therefore, the AGPs FermiNet provides a more powerful Ansatz with fewer variational parameters than the Slater FermiNet, since the former contains the latter.
It is worth mentioning another advantage of using the FermiNet to build geminals. By generating more sets of independent parameters $\mathbf{w}^{k}_{i}$ in Eq. (19), one can easily construct an arbitrary number $N_{\det}$ of FermiNet geminals $\varphi^{k}(\mathbf{r}^{\alpha}_{i},\mathbf{r}^{\bar{\alpha}}_{j};\{\mathbf{r}^{\alpha}_{/i}\};\{\mathbf{r}^{\bar{\alpha}}_{/j}\})$ with $k\in\{1,2,\dots,N_{\det}\}$, all without the use of a basis set. This allows one to use weighted sums of AGPs determinants as trial wave functions, similar to the weighted sum of conventional FermiNet determinants seen in Eq. (11).